5-13 Time for a Lump Sum to Double How long will it take $300 to double...

5-13

Time for a Lump Sum to Double How long will it take $300 to double if it earns the following rates? Compounding occurs once a year.

  1. 6%

  2. 13%

  3. 21%

  4. 100%
Answer #1

Rate Time 6% 11.90 5.67 3.64 13% 21% 100% 31

Rate Time 1 0.06 =NPER(B2,,-300,600) 2 0.13 =NPER(B3,,-300,600) 3 0.21 =NPER(B4,,-300,600) 41 =NPER(B5,,-300,600)